a. Find the sum of the series to n-lterms 2 1+Vx + 2 2 1- x 1-7X + + ... to n-l terms b. If the fifth term of the sequen
Answer #1

a Find the sum of series to bnl terms. 21+ +- to ni terms + 1-In It In а, a, a2 = 1-Jn Ito 1- 2 Q a₂-a, = I-n 1+. In (4-4-5)